# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control*

### v3

@@ Abstract
Practical control systems pose significant challenges in identifying optimal control policies due to uncertainties in the system model and external disturbances. While $H_{\infty}$ control techniques are commonly used to design robust controllers that mitigate the effects of disturbances, these methods often require complex and computationally intensive calculations. To address this issue, this paper proposes a reinforcement learning algorithm called Robust Deterministic Policy Gradient (RDPG), which formulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game. In this formulation, one player (the user) aims to minimize the cost, while the other player (the adversary) seeks to maximize it. We then employ deterministic policy gradient (DPG) and its deep reinforcement learning counterpart to train a robust control policy with effective disturbance attenuation. In particular, for practical implementation, we introduce an algorithm called robust deep deterministic policy gradient (RDDPG), which employs a deep neural network architecture and integrates techniques from the twin-delayed deep deterministic policy gradient (TD3) to enhance stability and learning efficiency.
To evaluate the proposed algorithm, we implement it on an unmanned aerial vehicle (UAV) tasked with following a predefined path in a disturbance-prone environment. The experimental results demonstrate that the proposed method outperforms other control approaches in terms of robustness against disturbances, enabling precise real-time tracking of moving targets even under severe disturbance conditions.
I INTRODUCTION
In recent years, significant advancements have been made in deep reinforcement learning (DRL), which integrates deep learning and reinforcement learning (RL). The successes of DRL have been demonstrated across various domains, including games [12, 13] and control [5, 14, 15]. Among the various challenging domains, this paper focuses on dynamic system control problems with continuous state-action spaces.
Practical dynamic control systems pose significant challenges in identifying optimal control policies due to unstructured uncertainties in the system model and external disturbances. Current state-of-the-art DRL methods for continuous control tasks, such as deep deterministic policy gradient (DDPG) [5], twin delayed DDPG (TD3) [4], soft actor-critic (SAC) [2], and proximal policy optimization (PPO) [3], generally fail to address these issues effectively. For instance, a DRL agent trained in a specific real-world or simulated environment may struggle to maintain optimal performance when deployed in different environments where the patterns of external disturbances changed. Therefore, developing a DRL agent policy that is robust to unseen disturbances is crucial for ensuring reliability in real-world applications.
The main goal of this paper is to introduce the robust deterministic policy gradient (RDPG) algorithm, which enhances the robustness of the DPG algorithm by integrating the concept of the $H_{\infty}$ control problem with a two-player zero-sum dynamic game framework.
The $H_{\infty}$ control method is a robust control technique that designs an optimal controller to minimize the impact of disturbances on its performance. It is well-known that, under certain assumptions, the $H_{\infty}$ control problem can be transformed into a two-player zero-sum dynamic game (or a minmax optimization problem), where the first player, or the controller, tries to minimize a specific cost, while the second player, or the adversary, attempts to maximize it [21]. The minmax optimization problem can be solved by the Hamilton–Jacobi–Isaacs (HJI) equation. If the system is linear with quadratic costs, the solution of HJI equation can be derived through a Riccati equation [22, 23]. However, if the system has nonlinearity or the cost is non-quadratic, the HJI equation is extremely difficult to solve directly. Therefore, some methods have tried to obtain the approximate solution of HJI equations including power series expansion-based method [6], locally linearized mode [7], online learning algorithm with neural nets [10], and relaxed policy iteration algorithm with bounded policy evaluation error [11]. However, these methods are still associated with a high computational cost.
@@ I INTRODUCTION
In this paper, we present the so-called robust deterministic policy gradient (RDPG) algorithm, which improves the robustness of DPG algorithm from the two-player zero-sum dynamic game perspectives. In particular, we transform the $H_{\infty}$ problem into a two-player zero-sum dynamic game between the user (controller) and the adversary that represents the disturbance. Then, we propose a version of deterministic policy gradient method to train both the user and adversary, where the user and adversary each learn their own policy to minimize and maximize the objective function, respectively. Consequently, the robust controller can determine the optimal control input, which is robust against disturbances. The proposed RDPG is then extended to a DDPG variant, called RDDPG (robust DDPG), which combines the idea of RDPG with TD3.
@@
Additionally, we aim to develop a robust control strategy for quadrotors, a commonly used unmanned aerial vehicle (UAV) across various applications. Ensuring robust quadrotor control is essential, particularly for tasks such as tracking moving targets or following predefined waypoints in dynamic environments. We demonstrate that the proposed RDDPG method effectively determines optimal control inputs in the presence of external disturbances. Furthermore, it outperforms other DRL-based control approaches by achieving lower cost under severe disturbance conditions.
@@
    We propose the robust deterministic policy gradient (RDPG) algorithm, which effectively improves the robustness against external disturbances by solving the $H_{\infty}$ problem through the dynamic game perspectives in combination with the actor-critic method.
@@
    Moreover, we develop a new DDPG algorithm, called RDDPG, which combines the concept of RDPG with TD3. Through numerical simulations of the quadrotor control problems, the effectiveness of RDDPG is demonstrated.
@@
    Through the developed method, we propose a robust tracking control design method for quadrotor applications. Comprehensive experimental results are given to demonstrate that the proposed approach significantly outperforms other state-of-the-art DRL approaches such as DDPG, PPG, SAC, and TD3.
II Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [6, 7, 8, 9]. However, these methods require precise model linearization and incur a high computational costs due to the need to solve HJI equations. To address this issue, alternative approaches leverage deep reinforcement learning (DRL) to determine the optimal control policy, which does not require solving the nonlinear HJI equation.
Until now, the two-player zero-sum game and $H_{\infty}$ control frameworks have been applied to deep reinforcement learning (DRL) to enhance robustness under various scenarios. For instance,[18] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. In [24], the authors proposed a novel approach for robust locomotion control of quadruped robots against external disturbances. A robust adversarial reinforcement learning strategy was proposed in [19] to address modeling errors and discrepancies between training and testing conditions. Although their underlying philosophy is similar to ours, their approaches are significantly different. The method in [24] solves two separate constrained optimization problems based on the ideas in PPO. The DRL in [18] primarily considers model uncertainties rather than external disturbances and applies the so-called action robust RL in [20], which has environmental structures different from our setting. The authors in [19] propose more general alternating optimization procedures and do not consider deterministic policy gradient.
III Background
III-A $H_{\infty}$ Control
Let us consider the discrete time nonlinear system
@@ III-A $H_{\infty}$ Control
Assume that the initial state $x_{0}$ is determined by $x_{0}\sim\rho(\cdot)$, where $\rho$ is the initial state distribution. Defining the stochastic processes ${\bf w}_{0:\infty}:=(w_{0},w_{1},\ldots)$ and ${\bf y}_{0:\infty}:=(y_{0},y_{1},\ldots)$, the system can be seen as a stochastic mapping from ${\bf w}_{0:\infty}$ to ${\bf y}_{0:\infty}$ as follows:
${\bf y}_{0:\infty}\sim T_{\pi}(\cdot|{\bf w}_{0:\infty}).$
where $T_{\pi}$ is the conditional probability of ${\bf y}_{0:\infty}$ given ${\bf w}_{0:\infty}$. Moreover, defining the $L^{2}$ norm for the general stochastic process ${\bf z}_{0:\infty}:=(z_{0},z_{1},\ldots)$ by
$||{{\bf{z}}_{0:\infty}}|{|_{{L^{2}}}}:=\sqrt{\sum\limits_{k=0}^{\infty}{\mathbb{E}}[z_{k}^{T}z_{k}]},$
@@
$||{T_{\pi}}|{|_{\infty}}:=\mathop{\sup}\limits_{{{\bf{w}}_{0:\infty}}\neq 0}\frac{{{{\left\|{{{\bf{y}}_{0:\infty}}}\right\|}_{{L^{2}}}}}}{{{{\left\|{{{\bf{w}}_{0:\infty}}}\right\|}_{{L^{2}}}}}}$ (2)
The goal of $H_{\infty}$ control is to design a control policy $\pi$ that minimizes the $H_{\infty}$ norm of the system $T_{\pi}$. However, minimizing $||T_{\pi}||_{\infty}$ directly is often difficult in practical implementations. Instead, we typically aim to find an acceptable $\eta>0$ and a controller satisfying $||T_{\pi}||_{\infty}\leq\eta$, which is called suboptimal $H_{\infty}$ control problem. Then, the $H_{\infty}$ control problem can be approximated to the problem of finding a controller $\pi$ that satisfies the constraint
$||{T_{\pi}}||_{\infty}^{2}=\mathop{\sup}\limits_{{{\bf{w}}_{0:\infty}}\neq 0}\frac{{||{{\bf{y}}_{0:\infty}}||_{{L^{2}}}^{2}}}{{||{{\bf{w}}_{0:\infty}}||_{{L^{2}}}^{2}}}\leq{\eta^{2}}$ (3)
@@
${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(y_{k}^{T}{y_{k}}-{\eta^{2}}w_{k}^{T}{w_{k}})}}\right|\pi}\right]$
@@
III-B Two-player zero-sum dynamic game perspective
According to [21], the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a zero-sum dynamic game under certain assumptions. In particular, we consider two decision making agents called the user and the adversary, respectively. They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form ${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$, where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the first agent, and $w_{k}$ is the control input of the second agent. The user’s primary goal entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs. There exist two categories of the two-player dynamic games, the alternating two-player dynamic game and simultaneous two-player dynamic game. In the alternating two-player dynamic game, two agents engaged in decision making take turns in selecting actions to maximize and minimize the cost, respectively. On the other hand, in the simultaneous two-player dynamic game, the two agents take actions simultaneously to maximize and minimize the cost. In this paper, we mainly focus on the simultaneous two-player dynamic game.
We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$
@@ III-B Two-player zero-sum dynamic game perspective
$c(x,u,w)=g{(x,u,w)^{T}}g(x,u,w)-{\eta^{2}}{w^{T}}w$
where $v\sim p(\cdot)$ and $\gamma=1$. The corresponding cost function is then written as
${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(y_{k}^{T}{y_{k}}-{\eta^{2}}w_{k}^{T}{w_{k}})}}\right|\pi,\mu}\right]$
@@ III-C Bellman equations
${Q^{\pi,\mu}}(x,u,w)$
::=:= ${\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=z,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
@@
IV Proposed method
In this section, we will present the proposed method.
IV-A Robust deterministic policy gradient
@@
The saddle-point problem is then converted to
@@
$\tilde{c}(x,u,w)=g{(x,u,w,v)^{T}}g(x,u,w,v)$
In this way, we can consider the two-player zero-sum dynamic game as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance.
@@
$\frac{{\mathbb{E}}\left[{\left.\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}(x_{k},u_{k},w_{k})}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{{\gamma^{k}}}w_{k}^{T}{w_{k}}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
IV-B Implementation based on TD3 (RDDPG)
To implement RDPG for practical high-dimensional tasks, we will apply the techniques from twin-delayed deep deterministic policy gradient (TD3) in [4], which is called robust deep deterministic policy gradient (RDDPG). In particular, we introduce the following networks:

### v4

@@ Abstract
Practical control systems pose significant challenges in identifying optimal control policies due to uncertainties in the system model and external disturbances. While $H_{\infty}$ control techniques are commonly used to design robust controllers that mitigate the effects of disturbances, these methods often require complex and computationally intensive calculations. To address this issue, this paper proposes a reinforcement learning algorithm called Robust Deterministic Policy Gradient (RDPG), which formulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game. In this formulation, one player (the user) aims to minimize the cost, while the other player (the adversary) seeks to maximize it. We then employ deterministic policy gradient (DPG) and its deep reinforcement learning counterpart to train a robust control policy with effective disturbance attenuation. In particular, for practical implementation, we introduce an algorithm called robust deep deterministic policy gradient (RDDPG), which employs a deep neural network architecture and integrates techniques from the twin-delayed deep deterministic policy gradient (TD3) to enhance stability and learning efficiency. To evaluate the proposed algorithm, we implement it on an unmanned aerial vehicle (UAV) tasked with following a predefined path in a disturbance-prone environment. The experimental results demonstrate that the proposed method outperforms other control approaches in terms of robustness against disturbances, enabling precise real-time tracking of moving targets even under severe disturbance conditions.
keywords:
Deep reinforcement learning, Two players zero-sum game, Unmanned aerial vehicles
1 Introduction
In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal policies across various domains, including games [1, 2] and control tasks [3, 4]. In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models. However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment. Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications. To address this challenge, in this paper, we introduce adversarial learning, where the agent is trained in the presence of disturbances. By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of DPG algorithm from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives. The $H_{\infty}$ control method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance. A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem. In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [8].
@@ 1 Introduction
Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between a user agent (controller) and an adversarial agent (disturbance generator). We then propose a reinforcement learning approach based on the deterministic policy gradient (DPG) method to simultaneously train both agents. Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it. As a result of this zero-sum game framework, the user learns a control policy that is inherently robust against worst-case disturbances. Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework. The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3.
@@
Additionally, we aim to develop a robust control strategy for quadrotors, a widely used unmanned aerial vehicle (UAV) in various applications. The ability to maintain robust control of quadrotors is particularly important in dynamic and uncertain environments, where tasks such as tracking moving targets or following predefined waypoints are required. We demonstrate that the proposed RDDPG method effectively determines optimal control inputs in the presence of external disturbances. Furthermore, it outperforms existing DRL-based control approaches by achieving lower cost under severe disturbance conditions.
@@
    We propose the robust deterministic policy gradient (RDPG) algorithm, which enhances robustness against external disturbances by integrating concepts from $H_{\infty}$ control and two-player zero-sum dynamic games into the actor-critic framework.
@@
    We develop a novel deep reinforcement learning algorithm, named RDDPG, which integrates RDPG into the TD3 framework to enable scalable and robust learning in continuous control tasks.
@@
    We apply the proposed RDDPG algorithm to the robust tracking control of quadrotors and conduct comprehensive numerical simulations. The results demonstrate that RDDPG significantly outperforms several state-of-the-art DRL methods, including DDPG, SAC, and TD3.
2 Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [9, 10, 11]. However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations. To address these limitations, alternative approaches leverage deep reinforcement learning (DRL) to determine optimal control policies without the need for solving nonlinear HJI equations.
Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios. For instance, [12] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [13], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward. In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary. By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise. In [14], the authors proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator. The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning. To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) [15] was proposed. RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system. The agent and the adversary are concurrently trained using the trust region policy optimization (TRPO) algorithm, with the agent learning to maximize the expected reward while the adversary learns to minimize it. Although the underlying concept is similar, their approach differs significantly from ours. Both methods adopt stochastic policy frameworks, such as PPO and TRPO, and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training. In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
3 Preliminaries
3.1 $H_{\infty}$ Control
Let us consider the discrete time nonlinear discrete-time system
@@ 3.1 $H_{\infty}$ Control
Assume that the initial state $x_{0}$ is determined by $x_{0}\sim\rho(\cdot)$, where $\rho$ is the initial state distribution. Defining the stochastic processes ${\bf w}_{0:\infty}:=(w_{0},w_{1},\ldots)$ and ${\bf y}_{0:\infty}:=(y_{0},y_{1},\ldots)$, the system can be seen as a stochastic mapping from ${\bf w}_{0:\infty}$ to ${\bf y}_{0:\infty}$ as follows: ${\bf y}_{0:\infty}\sim T_{\pi}(\cdot|{\bf w}_{0:\infty})$, where $T_{\pi}$ is the conditional probability of ${\bf y}_{0:\infty}$ given ${\bf w}_{0:\infty}$. Moreover, defining the $L^{2}$ norm for the general stochastic process ${\bf z}_{0:\infty}:=(z_{0},z_{1},\ldots)$ by
[…]
@@
[…]
The goal of $H_{\infty}$ control is to design a control policy $\pi$ that minimizes the $H_{\infty}$ norm of the system $T_{\pi}$. However, minimizing $T_{\pi}||_{\infty}$ directly is often difficult in practical implementations. Instead, we typically aim to find an acceptable $\eta>0$ and a controller satisfying $T_{\pi}||_{\infty}\leq\eta$, which is called suboptimal $H_{\infty}$ control problem. Then, the $H_{\infty}$ control problem can be approximated to the problem of finding a controller $\pi$ that satisfies the constraint
[…]
@@
${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi}\right]$
@@
3.2 Two-player zero-sum dynamic game perspective
According to [8], the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a zero-sum dynamic game under certain assumptions. In particular, we consider two decision making agents called the user and the adversary, respectively. They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form
${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$
where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the user, and $w_{k}$ is the control input of the adversary. The two agents have opposing objectives. The user’s primary objective entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs. There exist two categories of the two-player dynamic games, the alternating two-player dynamic game and simultaneous two-player dynamic game. In the alternating two-player dynamic game, two agents engaged in decision making take turns in selecting actions to maximize and minimize the cost, respectively. On the other hand, in the simultaneous two-player dynamic game, the two agents take actions simultaneously to maximize and minimize the cost. In this paper, we mainly focus on the simultaneous two-player dynamic game.
We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$ for the adversary
@@ 3.2 Two-player zero-sum dynamic game perspective
$c(x,u,w)=\left\|{g(x,u,w)}\right\|_{2}^{2}-{\eta^{2}}\left\|w\right\|_{2}^{2}.$
Then the corresponding cost function is then written when $\gamma=1$ as
${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi,\mu}\right]$
@@ 3.3 Bellman equations
${Q^{\pi,\mu}}(x,u,w):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=x,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
@@
4 Proposed method
In this section, we will present the proposed method, robust deterministic policy gradient (RDPG) and its deep reinforcement learning version, robust deep deterministic policy gradient (RDDPG).
Figure  1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game. In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment. The user policy generates the action $u_{t}$, while the adversary policy generates the disturbance $w_{t}$, both based on the current state $x_{t}$. These inputs are fed into the environment, which returns the next state $x_{t+1}$ and the cost $c_{t}$ based on the dynamics of the environment. The critic is updated based on $\{x_{t},u_{t},w_{t},x_{t+1},c_{t}\}$ and evaluates the joint performance of the user and adversary. The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi,\mu}$. The detailed procedure is described in the following sections.
Figure 1: Overview of RDDPG. Based on the two-player zero-sum dynamic game, the user policy $\pi$ seeks to minimize the objective function $J^{\pi,\mu}$, while the adversary policy $\mu$ attempts to maximize it by injecting disturbances into the environment.
4.1 Robust deterministic policy gradient
@@
where $\pi_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denotes the user’s control policy parameterized by $\theta$ and $\mu_{\phi}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ denotes the adversary’s control policy parameterized by $\phi$.. The saddle-point problem is then converted to
@@
$\tilde{c}(x,u,w)=\left\|{g(x,u,w,v)}\right\|_{2}^{2}$
In this way, we can consider the two-player zero-sum dynamic game as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance. The additional disturbance quadratic cost ${\eta^{2}}w^{T}w$ in the cost function discourages the adversary from generating excessively large disturbances, effectively regularizing its behavior and ensuring that the learned policies are robust against realistic worst-case scenarios rather than arbitrary or unbounded inputs.
@@
$\frac{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}({x_{k}},{u_{k}},{w_{k}})}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}\gamma^{k}\left\|w_{k}\right\|_{2}^{2}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
4.2 Robust deep deterministic policy gradient
To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the robust deep deterministic policy gradient (RDDPG). In particular, we introduce the following networks: